OpenStudy (anonymous):

If h(x) = −4x − 10, find h(−5). −30 1.25 10 −1

OpenStudy (nurali):

h(x) = −4x − 10 h(−5)=-4(-5)-10 h(-5)=20-10 h(-5)=10

OpenStudy (ivettef365):

all you need to do is substitute the x on function h(x) for -5